# Scalable magnetic resonance fingerprinting: Incremental inference of high dimensional elliptical mixtures from large data volumes

### v1

@@ 7.5 Computation times
Experiments conducted on an Nvidia V100-32gb GPU are fast. It takes only $7$ minutes to reconstruct $1$ slice of a subject using both HD-GMM and HD-STM, compared to $45$ minutes for full matching. This represents a $6$-fold time reduction for an HD-MED model. However, in more realistic scenarios, calculations are performed locally by a medical practitioner on a CPU. On an Apple M2 Pro CPU, full matching takes days and is prone to many memory issues that needs to be solve, while the HD-GMM and HD-STM variants require only $304h$ 30.
@@ 8 Conclusion
In this work, we combine robust latent variable representations, clustering and incremental learning to propose a new tractable and accurate way to represent and handle large volumes of potentially heterogeneous high-dimensional data. To our knowledge, this combination and its use as a data compression strategy is novel. The clustering structure of HD-MED allows to address data heterogeneity, for a greater dimensionality reduction without increasing information loss. Incremental learning allows to handle large volumes, resulting in significant reduction in both computational costs and information losses. In terms of implementation, the procedure is flexible and easy to interpret. It depends on two main hyperparameters, the number of clusters $K$ and the vector of reduced dimensions ${\mathb{d}}$, that can be set using conventional model selection criteria or chosen by the user to meet resource constraints. Typically increasing $K$ to reduce the size of clusters or decreasing ${\mathb{d}}$ to increase compression.
@@
As an illustration, we focus on a Magnetic Resonance Fingerprinting (MRF) application. The proposed method drastically reduces the computational time required on standard hardware, such as CPUs, which are commonly used in clinical environments. Beyond this crucial processing time gain, dimensionality reduction has also an interesting impact on patient data acquisition time, as it helps mitigating the effect of noise in fast-acquired in vivo signals. This reductions are thus not only a technical achievement but also a significant step forward in making advanced MRF techniques more accessible and usable in everyday clinical practice.
@@
In addition, although illustrated with a simple matching procedure, which is the current reference in the target MRF application, our procedure can be coupled with other simulation-based inference (SBI) approaches (Cranmer, Brehmer and Louppe, 2020), such as approximate Bayesian computation (ABC) or other neural techniques (Barrier et al., 2024). Like MRF, simulation-based inference has to face two opposite requirements, which are the need for large and high-dimensional simulated data sets to accurately capture information on the physics under study and the issue of handling such large volumes due to computational ressources constraints. ABC uses distances between observations and simulations, and inference is based on these distances. A reduced data representation can be typically coupled with ABC techniques proposing automatic summary statistics selection such as in Forbes et al. (2022); Fearnhead and Prangle (2012). It can also be used with SBI neural approaches, e.g. Häggström et al. (2024), when coupled with a regression model for high-dimensional data, see Algorithm 1 in supplementary material. Compare to matching and standard ABC, regression-based procedures are more amortized, which further reduces the reconstruction time. In terms of MRF and other medical imaging applications, future work will thus involve combining HD-MED with a regression model or a neural network, such as in Deleforge, Forbes and Horaud (2015) or Golbabaee et al. (2019).
@@
At last, none of the previously mentioned methods, currently make use of spatial information at the voxel level, to perform map reconstruction. The idea would be to exploit voxel proximity to either improve or accelerate parameter prediction. A relatively easy extension of mixture-based methods is to account for spatial information by adding a Markov dependence on the clusters, as done for instance in Deleforge et al. (2015).
9 Financial disclosure
G. Oudoumanessah was financially supported by the AURA region, and was granted access to the HPC resources of IDRIS under the allocation 2022-AD011013867 made by GENCI.
@@ 9 Financial disclosure
The MRI facility IRMaGe is partly funded by the French program “Investissement d’avenir” run by the French National Research Agency, grant “Infrastructure d’avenir en Biologie et Santé” [ANR-11-INBS-0006]. The project is supported by the French National Research Agency [ANR-20-CE19-0030 MRFUSE].

### v2

@@ 7.5 Computation times
Training is performed on an Nvidia H100-80GB GPU, requiring 26 hours for HD-GMM, 28 hours for HD-STM, and 12 hours for DRONE. For HD models, these times do not include initialization, which consists of computing BIC for different $K$ on a subset of signals. The different $K$ values can be run in parallel so that initialization time mainly depends on the maximum $K$ tested and the number of sub-sampled signals. As an illustration, for $K=80$ and $10,000$ signals, initialization takes 15 minutes. Reconstruction is executed on an Nvidia V100-32GB GPU, taking 3 minutes $\pm$ 2 seconds per slice for both HD-GMM and HD-STM, significantly faster than the 29 minutes $\pm$ 46 seconds required for full matching. Due to memory constraints, as the full uncompressed dictionary exceeds the 32GB GPU memory capacity, an incremental matching approach was implemented. Slice reconstruction using DRONE on an Nvidia V100-32GB GPU is notably rapid, completing in under 1 minute. This highlights the efficiency of neural network-based regression methods. However, DRONE exhibits limitations in accurately estimating vascular parameters (Figure 6). In more realistic scenarios, calculations are performed locally by a medical practitioner on a CPU. On an Apple M2 Pro CPU, full matching takes days and is prone to many memory issues that need to be solved, while the HD-GMM and HD-STM variants require only $4.5$ hours. This is a strength of HD models over complex neural network approaches, which cannot be easily adapted to CPU and whose running times would be prohibitive.
@@ 8 Conclusion
In this work, we combine robust latent variable representations, clustering and incremental learning to propose a tractable and accurate way to represent and handle large volumes of potentially heterogeneous high-dimensional data. To our knowledge, this combination and its use as a data compression strategy is novel. The clustering structure of HD-MED allows to address data heterogeneity, for a greater dimensionality reduction without increasing information loss. Incremental learning allows to handle large volumes, resulting in significant reduction in both computational costs and information losses. In terms of implementation, the procedure is flexible and easy to interpret. It depends on two main hyperparameters, the number of clusters $K$ and the vector of reduced dimensions $\mathb{d}$}, that can be set using conventional model selection criteria or chosen by the user to meet resource constraints, e.g. increasing $K$ to reduce the size of clusters or decreasing $\mathb{d}$} to increase compression.
@@
As an illustration, we focus on a Magnetic Resonance Fingerprinting (MRF) application. The proposed method drastically reduces the computational time required on standard hardware, such as CPUs, which are commonly used in clinical environments. Beyond this crucial processing time gain, dimensionality reduction has also an interesting impact on patient data acquisition time, as it helps mitigate the effect of noise in fast-acquired in vivo signals. This reduction is thus not only a technical achievement but also a significant step forward in making advanced MRF techniques more accessible and usable in daily clinical practice. Currently, the sequence used to acquire these images is in the pre-clinical testing phase on pathological animal, including those with tumors and stroke. With respect to cross-vendor data, since MRF provides quantitative measurements, there are fewer harmonization issues between scanners from different vendors. However, a potential limitation is the noise, which can differ between scanners; our simulated dictionary originally contains denoised signals, which we artificially noised with different values of the signal-to-noise ratio (SNR), but we did not adjust the noise to exactly match that of the acquisition data, also suggesting a potentially good generalization between acquisitions from different scanners.
@@
In addition, although illustrated with a simple matching procedure, which is the current reference in the target MRF application, our procedure can be coupled with other simulation-based inference (SBI) approaches (Cranmer), such as approximate Bayesian computation (ABC) or other neural techniques (barrier2024). Like MRF, simulation-based inference has to face two opposite requirements, which are the need for large and high-dimensional simulated data sets to accurately capture information on the physics under study and the issue of handling such large volumes due to computational resource constraints. ABC uses distances between observations and simulations, and inference is based on these distances. A reduced data representation can typically be coupled with ABC techniques proposing automatic summary statistics selection such as in forbes2022; fearnhead2012. It can also be used with SBI neural approaches, e.g. haggstrom2024, when coupled with a regression model for high-dimensional data, see Algorithm 1 in supplementary material. Compared to matching and standard ABC, regression-based procedures are more amortized, which further reduces the reconstruction time. In terms of MRF and other medical imaging applications, future work will thus involve combining HD-MED with a regression model or a neural network, such as in Deleforge2015 or golbabaee2019. More generally, the possibility to incrementally learn HD-MED models, from high-dimensional heterogeneous data in large volumes, opens the way to a broad range of applications. A first type of applications relates to anomaly detection using a likelihood score, see Lagogiannis2024 for a recent review. Anomaly detection can be described as an outlier detection problem, where the aim is to discriminate between in- and out-of-distribution samples of a normative distribution learned from a dataset of normal samples. These normal samples can typically be high-dimensional feature vectors summarizing a possibly heterogeneous set of subjects. Second, inverse problems in an SBI context provide numerous applications other than MRF, such as remote sensing tasks in planetary science, extending, for instance, the work by Kugler2022. Third, HD-MED could also simply be used as an alternative to standard PCA to achieve a better low-dimensional representation of datasets of interest, or simply as a clustering technique (bouveyron2014).
@@
Lastly, none of the mentioned methods makes use of spatial information, to perform map reconstruction. The idea would be to account for voxel proximity to improve or accelerate prediction. A natural extension of mixture-based methods is to account for spatial information by adding a Markov dependence on the clusters, as done by DeleforgeFBH15.
{acks}
[Acknowledgments] The authors would like to thank the referees, an Associate Editor and the Editor for their constructive comments that improved the quality of this paper.
{funding}
@@
[Financial disclosure] GO was supported by the AURA region and was granted access to the HPC resources of IDRIS under the allocation 2022-AD011013867 made by GENCI. The MRI facility IRMaGe is partly funded by the program “Investissement d’avenir” run by the French National Research Agency, grant “Infrastructure d’avenir en Biologie et Santé” [ANR-11-INBS-0006]. The project is supported by the French National Research Agency [ANR-20-CE19-0030 MRFUSE].
{supplement}
Supplementary material includes an additional discussion on the pratical benefits of the online EM algorithm. Details on its application to finite mixture models, and in particular to mixtures of Student and Laplace distributions in a high-dimensional settting are also provided. More illustrations on the fingerprinting application can also be found. The code to run the proposed algorithm and reproduce key results is also accessible on GitHub (https://github.com/geoffroyO/onlineHDEM). The repository includes a notebook, tutorial_hd_models.ipynb, which provides a comprehensive tutorial on how to run the different models with simulated data. However, for regulatory and patient privacy reasons, the data used in this study cannot be made publicly available.